TWiki home Point 1 Point 2 Point 3 Point 4 Point 5 Point 6 Point 7 Point 8 Minutes


Created by Main.MircoCoccoli on 2006-11-28 - 16:40

Preparation for Cool Down sec. 78 - Nov 28

PDF RAT_20061128.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2006x11x28x16x37/RAT_20061128.pdf

-- Main.MircoCoccoli - 28 Nov 2006 - 16:38

Open Issues:

AC non-conformity

  • 21.11.06 Non-conformity on instrumentation cables (temperature sensors on current leads) (A.Suraci: 4 cables are in repair-Wed.29th)
  • 13.11.06 Non-conformity of the crates of cryo instrumentation (inrush current) (A.Suraci: ready next year)
  • 13.11.06 Non-conformity of the AC cabling of the crates under the magnets. This concerns ACR, MEL, orbit corrector power converters & VAC.
DC cabling
  • 25.11.06 Functional test on the leads heaters: DFBAO, DFBMA, DFBMC, DFBXG, DFBMH
  • 22.11.06 Pre-connection of 120A cables in LSSR7 (TS/EL)
  • 20.11.06 Galvanic insulation installation on all DFBs (AT/MEL)
  • 13.11.06 Non-conformity of the DC cabling of the orbit corrector power converters
  • 13.11.06 6 kA and 13 kA cables positioning at DFBAO, DFBMA, DFBMC, DFBAN (Flohe)
Assembly - Vacuum - Cryogenics
  • 22.11.06 Cryo-valves remote control in CCC by AB/CO
  • 22.11.06 Control of quench valve
  • 20.11.06 DFB instrumentation cables to be connected (A.Suraci: 80% done)
  • 17.11.06 Instrumentation cable HV tests in LSS8L (A.Suraci: 4 cables to be tested)
DFB Commissioning
  • 13.11.06 ElQA-TP4-A for all DFBs between 7 and 8 (D.Bozzini: (done: DFBAO; next(w48): DFBAN, DFBMA, DFBMC, DFBMH, DFBX))
  • 17.11.06 Pumping of insulation vacuum for all DFBs between 7 and 8 (P.Chambouvet: DFBAN done; next: DFBMA, DFBMC, DFBX)
Safety General remark: written communication in advance to announce operations!!!
  • 25.11.06 ElQA-TP4B safety procedure (1900 volts tests!).
  • 23.11.06 Wooden structure for UA access restriction (that will allow transport).
  • 23.11.06 Preparation of a prototype insulating protection (derived from a plastic tube) of the monorail by J-C.Guillaume and approval by J.Etheridge.
  • 23.11.06 Water leak on the tunnel concrete wall to be fixed (33L8).
  • 16.11.06 Cool down safety procedure & access conditions
Preparation of powering test
  • 22.11.06 Pretest of 60A converters in sector 78, location C16L8. (B.Bellesia: stopped due to LQA hardware problems; expected end by Tuesday 28 Nov)
  • 22.11.06 Installation of flexible cables by CV (UA83)
  • 16.11.06 Water circulating in UA83 for the Powering Procedure Test: filters need to be checked.
AOB
  • 13.11.06 BPM connection in Q2, Q3, Q4, Q5, Q6, Q7–wk49 (J-C.Guillaume & R.Jones)
  • 13.11.06 Quench protection continuity tests and cables assignment LLS7R (DFBMH, DFBAN) and LSS8L (DFBMC, DFBMA and DFBX)

Edit, attach file or add comment

Edit | Attach | Watch | Print version | History: r1 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r1 - 2006-11-01 - JacekSzkutnik
 
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    HCC All webs login